eli designed a rectangular garden with length of 5 meters less than twice its width. if the perimeter of the garden is 50 meters, find the dimensions of the garden.
Answer by waynest(281)

length = 2w - 5
width = w
perimeter = 50
2(2w - 5) + 2w = 50
4w - 10 + 2w = 50
6w - 10 = 50
6w - 10 + 10 = 50 + 10
6w = 60
6w/6 = 60/6
w = 10
check:
2(2(10) - 5) + 2(10) = 50
2(20 - 5) + 20 = 50
2(15) + 20 = 50
30 + 20 = 50
50 = 50